OpenStudy (anonymous):

Find the sum of an 8-term geometric sequence when the first term is 7 and the last term is 114,688 and select the correct answer below. 152,915 16,384 16,377 152,908

OpenStudy (solomonzelman):

have you seen this: \(\large\color{black}{ \displaystyle a_n=a_1\times{\rm r}^{n-1} }\) ?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle a_n=a_1\times{\rm r}^{n-1} }\) your number of terms is 8 \(\large\color{black}{ \displaystyle a_8=a_1\times{\rm r}^{8-1} }\) for 8th term. \(\large\color{black}{ \displaystyle a_8=a_1\times{\rm r}^{7} }\) now plug in 8th term and 1st term to find the common ratio. \(\large\color{black}{ \displaystyle 114688=7\times{\rm r}^{7} }\)

OpenStudy (solomonzelman):

can you fsolve for common ratio (r) ?

OpenStudy (solomonzelman):

don't be silent.... because i can't "talk to a mute".

OpenStudy (solomonzelman):

with a mute*